OpenStudy (anonymous):

What are the solutions of 3x2 + 15 = -6x? -1 ± i -1 ± 2i -2 ± i -2 ± 2i

OpenStudy (anonymous):

okay let's solve it let's 3x^2+6x+15=0 let take three as common factor we get 3(x^2+2x+5)=0 now as \[3\neq0\] we have x^2+2x+5=0 by comparing it with standard form ax^2+bx+c we get a=1 b=2 c=5 but this values in \[x=-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

 Would it look like this?

OpenStudy (anonymous):

correct ..! genius..! now let's further simplify it.. we can write by simplification as \[x=-2 \pm (\sqrt{16} \sqrt{-1})/2\] as you know \[\sqrt{-1}=i\] and \[\sqrt{16}=4\] so we can write \[x=-2\pm 4i/2\] which is equal to \[x=-1\pm2i\] i hope you got it.!! :)

OpenStudy (anonymous):

Thanks. :) Can you help me with one more?

OpenStudy (anonymous):

yes i would love to..! let's solve it

OpenStudy (anonymous):

What are the solutions of 4x2 + x = -3? the quantity negative 1 plus or minus 7i all over 8 the quantity negative 1 plus or minus i square root of 5 all over 2 the quantity negative 1 plus or minus i square root of 47 all over 8 the quantity negative 1 plus or minus 2i square root of 7 all over 8

OpenStudy (anonymous):

I've gotten this so far: 

OpenStudy (ranga):

4x^2 + x = -3 4x^2 + x + 3 = 0 a=4, b=1, c=3 Substitute a, b and c in the quadratic formula.

OpenStudy (ranga):

\[x = \frac{ -1 \pm \sqrt{(-1)^{2} - 4(4)(3)} }{ 2(4) }\]

OpenStudy (ranga):

\[x = \frac{ -1 \pm \sqrt{1 - 48} }{ 8 } = \frac{ -1 \pm i \sqrt{47} }{ 8 }\] Which answer matches the above?